# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ 7 Conclusion
We proposed a general approach called CRS to optimize noise schedules for training and sampling in diffusion models. Through extensive experiments, we confirmed that CRS broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and NFEs ranging from 5 to 250. CRS is fundamentally different from conventional approaches noise_ays ; noise_gits in that it can optimize the training schedule as well and use any distance metric for measuring probability-distributional change, allowing the introduction of domain specific distance metric and the use of different noise schedules for training and sampling. Although this study has the limitations discussed in Sec. 6, we believe that it successfully provides solid baselines. We hope that future research will further improve the generality of CRS, leading to its widespread adoption as a general-purpose approach.
@@ Appendix A Probability Distribution of Diffused Data
We derive the probability distribution of diffused data used in the toy example (Fig. 1).
The probability distribution of diffused data is given by
[…]
@@
where $q(\bm{x}_{0})$ is the probability distribution of a target dataset. We then approximate $q(\bm{x}_{0})$ with the empirical distribution as
@@
where $N$ is the number of samples in the target dataset, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes Dirac’s delta function. By substituting Eq. (19) into Eq. (17), we obtain
@@
We used the following formula of Dirac’s delta function:
[…]
@@
Appendix B Derivation of Distance Metric Based on Data Prediction from KLUB
We show that $v_{x}(\alpha)$ can be derived as a simplified version of KLUB introduced in AYS.
@@ B.1 KLUB in Variance-Preserving Process
The forward SDE of the continuous-time diffusion models [19] is defined as
@@
where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The associated reverse SDE is given by
@@
For approximately computing the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train the following noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
@@
The noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ has the following relationship with the data prediction model $\bm{x}_{\theta}(\bm{x},\alpha)$:
@@
Stochastic DDIM is a stochastic sampler designed to numerically solve the reverse SDE described above. In Stochastic DDIM, the solution for the first term of $\bm{f}(\bm{x},t)$ (linear term) is exactly computed, and discretization errors arise only from the second term of $\bm{f}(\bm{x},t)$ [22, 23]. According to a previous study [16], in the sub-interval $[t_{i-1},t_{i}]$, Stochastic DDIM solves the following discretized SDE:
@@
The KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by exactly solving the reverse SDE and approximately solving the reverse SDE using Stochastic DDIM is introduced as
[…]
The diffusion coefficient $g(t)$ in the variance-preserving process is given by
@@
By substituting Eq. (32) into Eq. (31), we obtain
[…]
B.2 KLUB as Distance Metric of CRS
We transform the KLUB into a form that can be applied to CRS. Because the KLUB is used to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
First, we approximate the integral in the KLUB using Simpson’s rule as follows:
@@ B.2 KLUB as Distance Metric of CRS
Next, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (36)
@@
$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0}),$ (37)
@@
We use $\overline{D_{x}^{2}}(\alpha,\alpha)=0$ in the second equality. Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the distance metric $D(t,t^{\prime})$ in CRS, we obtain the following equation:
@@
which is the same form as Eq. (7). We can regard $v_{x}(\alpha)$ as a simplified version of $v_{\mathrm{KLUB}}(\alpha)$ in the sense that the weighting function of $v_{\mathrm{KLUB}}(\alpha)$, namely $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$, is discarded in $v_{x}(\alpha)$.
It is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$. However, when $v_{\mathrm{KLUB}}(\alpha)$ is directly used, the noise schedule is heavily impacted by the weighting function rather than $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. The weighting function becomes extremely large around $\alpha\simeq 1(\sigma\simeq 0)$, which overwhelms the impact of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. Recall that $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ can be interpreted as measuring probability-distributional changes of diffused data, as described in Sec. 4.3. Therefore, the simplified version we introduce is more aligned with our objective of optimizing the noise schedule on the basis of distributional changes.
@@ Appendix C Implementation of CRS
We describe how to compute $v(\alpha)$ using each distance metric.
C.1 Distance Metric Based on FID
The forward process is simulated using the training dataset for computing $v_{\mathrm{FID}}(\alpha)$, as shown in Algorithm 1. One of the inputs $\{\alpha_{t}\}_{t=0,...,T}$ needs to be discretized to approximately represent the functional form of $v(\alpha)$ through linear interpolation. In our experiments, we set
@@ C.1 Distance Metric Based on FID
where $T=1000$, and set $p$ to $2$ and $1$ for the diffusion processes in pixel space and latent space, respectively. For the feature model $\bm{\phi}$, we used the Inception-V3 model trained on ImageNet classification for the pixel-space diffusion process. For the latent-space diffusion process, however, we need a feature model trained in the latent space embedded using the autoencoder. We trained ResNet50 on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model $\bm{\phi}$. Note that a trained diffusion model is not necessary for computing $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain noise schedules by solving Eq. (8).
The dependence of CRS-$v_{\mathrm{FID}}$ on hyperparameter $\xi$ is illustrated in Fig. 7. CRS-$v_{\mathrm{FID}}$ is computed on LSUN Horse 256 $\times$ 256 for the pixel-space diffusion process. As $\xi$ is increased, more timesteps are allocated to the region with low-noise level.
The computation time of $v_{\mathrm{FID}}(\alpha)$ is proportional to the number of images in the training dataset. For the pixel-space diffusion process, it takes about 35 hours for LSUN Horse 256 $\times$ 256 with 2M images using eight A100 GPUs. For the latent-space diffusion process, it takes about 12 hours for ImageNet 256 $\times$ 256 with 1.28M images using eight A100 GPUs. The computation of $v_{\mathrm{FID}}(\alpha)$ needs to be executed only once before training the diffusion model, and the computation time is sufficiently small compared with the training time of the diffusion model.
@@
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, number of timesteps for simulating the forward process $T$, discretized noise levels $\{\alpha_{t}\}_{t=0,...,T}$, feature model $\bm{\phi}$
@@
1:  # Compute statistics of feature vectors
2:  for $n=0$ to $N-1$ do
3:     $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
4:     if diffusion process in latent space then
5:        $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
6:     end if
7:
8:     # Simulate the forward process
9:     for $t=1$ to $T$ do
10:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
11:        $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
12:        $x^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
13:     end for
14:  end for
15:  $\bm{\mu}_{t}\leftarrow mean\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
16:  $\bm{\Sigma}_{t}\leftarrow cov\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
@@
18:  # Compute $v(\alpha)$ using $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$
19:  for $t=0$ to $T-1$ do
20:     $v_{t}\leftarrow\frac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
21:  end for
22:  $v_{T}\leftarrow v_{T-1}$
Figure 4: Dependence of CRS-$v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$.
Figure 5: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $w_{x}$ and $w_{\mathrm{FID}}$ ($\xi_{x}=\xi_{\mathrm{FID}}=1)$.
Figure 6: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
Figure 7: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{x}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
C.2 Distance Metric Based on Data Prediction
Algorithm 2 presents the pseudocode for computing $v_{x}(\alpha)$ using a trained diffusion model. This pseudocode focuses on the optimization of sampling schedules. Similar to the computation of $v_{\mathrm{FID}}(\alpha)$, the forward process is simulated using the training dataset, and the change in the data prediction between adjacent timesteps is computed. A common configuration of hyperparameters $(T=10^{3},S=10^{4},\alpha_{s}=1,\alpha_{e}=10^{-4})$ was used in our experiments.

### v4

@@ 7 Conclusion
We proposed a general framework called CRS for optimizing noise schedules in diffusion models. CRS enforces a constant rate of distributional change throughout the diffusion process and provides a unified perspective for designing both training and sampling schedules. Our framework supports arbitrary user-specified discrepancy measures, and we introduced three practical options that capture different aspects of the distributional dynamics.
Through extensive experiments, we demonstrated that CRS consistently improves the performance of both pixel-space and latent-space diffusion models across various datasets, samplers, and NFEs ranging from 5 to 250. In particular, for sampling in pixel-space diffusion models, CRS-$v_{x}+v_{\mathrm{FID}}$ with tuned hyperparameters achieved the best results. In all other settings, CRS-$v_{x}$ with $\xi=1$ performed consistently well and is recommended as a default configuration due to its simplicity and robustness.
Although this study has limitations, as discussed in Sec. 6—including the lack of a complete theoretical analysis and the need for further validation in other domains—we believe that CRS provides a solid and extensible foundation for future research on schedule optimization in diffusion models. We hope that this work encourages broader exploration of discrepancy-based scheduling, and that future research will further generalize CRS and extend its applicability beyond image generation.
Acknowledgments
We thank the anonymous reviewers for their constructive and insightful comments, which helped improve the quality of this paper. We are also grateful to the team responsible for operating and maintaining the AI Cloud Platform, which provides the internal GPU infrastructure used in this work. Without their support, this research would not have been possible.
@@ Appendix A Probability Distribution of Diffused Data
We derive the probability distribution of the diffused data used in the toy example (Fig. 1).
The distribution of the diffused data is given by:
$q(\bm{x}_{\alpha})$ $=\int q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})d\bm{x}_{0},$ (18)
@@
where $q(\bm{x}_{0})$ is the probability distribution of the target dataset.
We approximate $q(\bm{x}_{0})$ by the empirical distribution:
@@
where $N$ is the number of data samples, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes the Dirac delta function. Substituting Eq. (20) into Eq. (18) yields:
@@
This derivation uses the following formula of Dirac’s delta function:
$\int f(\bm{x})\delta(\bm{x}-\bm{y})d\bm{x}=f(\bm{y}).$ (22)
@@
Appendix B Derivation of Discrepancy Measure Based on Data Prediction from KLUB
We show that $v_{x}(\alpha)$ can be derived as a simplified variant of the KLUB introduced in AYS.
@@ B.1 KLUB in Variance-Preserving Process
The forward SDE in continuous-time diffusion models (Song et al., 2021b) is defined as:
@@
where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The corresponding reverse SDE is given by:
@@
To approximate the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train a noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
@@
The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is related to the noise prediction via:
@@
Stochastic DDIM numerically solves this reverse SDE. It exactly integrates the linear term and approximates the remaining term, introducing discretization error (Lu et al., 2022a; b). According to Sabour et al. (2024), in the interval $t_{i-1},t_{i}]$, Stochastic DDIM solves the discretized SDE:
@@
Furthermore, the KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by solving the true reverse SDE and the descretized one is introduced as
$\mathrm{KLUB}(t_{i-1},t_{t})$ $=\frac{1}{2}\int_{t_{i-1}}^{t_{i}}\mathbb{E}_{{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}}\left[\frac{\left\|\bm{f}(\bm{x}_{t},t)-\bm{f}^{\prime}(\bm{x}_{t},t)\right\|_{2}^{2}}{g(t)^{2}}\right]dt,$ (31)
$=\frac{1}{2}\int_{t_{i-1}}^{t_{i}}\mathbb{E}_{{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}}\left[\frac{\alpha(t)^{2}}{\sigma(t)^{4}}g(t)^{2}\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i}))\right\|_{2}^{2}\right]dt.$ (32)
In the variance-preserving setting, the diffusion coefficient satisfies:
@@
Substituting Eq. (33) into Eq. (32) gives:
$\mathrm{KLUB}(t_{i-1},t_{i})=-\int_{t_{i-1}}^{t_{i}}dt\dot{\alpha}(t)\frac{\alpha(t)}{\sigma(t)^{4}}\mathbb{E}_{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t_{i}}\sim q(\bm{x}_{t_{i}}|\bm{x}_{t})\end{subarray}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i}))\right\|_{2}^{2}\right].$ (34)
B.2 KLUB as Discrepancy Measure of CRS
To apply KLUB within the CRS framework, we reinterpret it as a discrepancy measure over adjacent steps in the diffusion process. Because the KLUB is introduced to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
We approximate the integral in the KLUB using Simpson’s rule as follows:
@@ B.2 KLUB as Discrepancy Measure of CRS
Then, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right],$ (37)
@@
$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=\int q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})d\bm{x}_{0},$ (38)
@@
We use $\overline{D_{x}^{2}}(\alpha,\alpha)=0$ in the second equality. Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the discrepancy measure $D(t,t^{\prime})$ in CRS, we obtain the following equation:
@@
which corresponds to the general form of CRS in Eq. (8). We can interpret $v_{x}(\alpha)$ as a simplified variant of $v_{\mathrm{KLUB}}(\alpha)$, in which the multiplicative weighting term $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$ is omitted.
Although it is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$ directly, this can lead to undesirable behavior, as the weighting term $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$ grows very large when $\alpha\simeq 1$ (i.e., $\sigma\simeq 0$), which can overwhelm the influence of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. Since $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ can be interpreted as capturing the distributional change between diffused data distributions (as discussed in Section 4.3), the simplified variant $v_{x}(\alpha)$ is more aligned with the objective of CRS.
@@ Appendix C Implementation of CRS
We describe how to compute $v(\alpha)$ using each discrepancy measure.
C.1 Discrepancy Measure Based on FID
Algorithm 1 presents the pseudocode for computing $v_{\mathrm{FID}}(\alpha)$, and hyperparameters are summarized in Tabel 8. Once the forward process is defined, $v(\alpha)$ can be computed by simulating the forward diffusion using the training dataset. Since this procedure is independent of model parameters, it only needs to be executed once prior to training the diffusion model.
It is important to note that, in latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded by the autoencoder, as the Gaussian approximation in Eq. (4) applies to the diffusion process defined in the latent space. Therefore, for latent-space models, we trained a ResNet-50 (He et al., 2016) classifier on ImageNet in the autoencoder’s latent space and use it as the feature extractor $\bm{\phi}$.
To compute $v(\alpha)$, we discretize the range of $\alpha$ using a one-dimensional grid $\{\alpha_{t}\}_{t=0,\dots,T}$ and approximate the continuous function via linear interpolation. In our experiments, we define the discretized grid of $\alpha$ values as:
@@ C.1 Discrepancy Measure Based on FID
where $T=1000$. For pixel-space diffusion models, where $v(\alpha)$ exhibits sharp variation near $\alpha\simeq 1$, we set $p=2$ to allocate more evaluation points in that region. For latent-space models, where $v(\alpha)$ is more uniformly distributed, we use $p=1$.
The computational cost of $v(\alpha)$ is proportional to the number of training images. For the pixel-space diffusion process on LSUN Horse $\times256$ with 2M images, the computation takes approximately 35 hours using eight A100 GPUs. For the latent-space process on ImageNet $\times256$ with 1.28M images, it takes about 12 hours using eight A100 GPUs. To address the scalability issue for large-scale datasets, we introduce a practical alternative in Sec. 5.7.
@@
1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N-1}$, timesteps $T$, noise levels $\{\alpha_{t}\}_{t=0,\dots,T}$, feature model $\bm{\phi}$
@@
3:# Compute feature vectors of diffused data
4:for n=0n=0toN1N-1 do
5:  $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
6:  if is_latent_space then
7:   $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
8:  end if
9:
10:  # Simulate the forward process
11:  for t=1t=1toTT do
12:   $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
13:   $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
14:   $\bm{x}^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\,\mathcal{N}(\bm{0},\bm{I})$
15:  end for
16:end for
@@
18:# Compute statistics of feature vectors
[…]
21:
22:# Compute $v(\alpha)$ based on FID
23:for t=0t=0toT1T-1 do
24:  $v_{t}\leftarrow\dfrac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
25:end for
[…]
Table 8: Hyperparameters for computing $v_{\mathrm{FID}}(\alpha)$.
Parameter Description Used Value
$T$
Number of timesteps
for simulating the forward process
$10^{3}$
$\{\alpha_{t}\}_{t=0,...,T}$
Grid of $\alpha$-values
for evaluating $v(\alpha)$
$\alpha_{t}=1-\left(\frac{t}{T}\right)^{p},$
where $p=\begin{cases}2,&\mathrm{pixel\ space}\\ 1,&\mathrm{latent\ space}\end{cases}$
$\phi$
Feature extractor
for computing FID
Pixel space: Inception-V3 model used
in the EDM repository for FID evaluation
Latent space: ResNet-50 trained by us
on ImageNet classification in the latent space
C.2 Discrepancy Measure Based on Data Prediction
Algorithm 2 presents the pseudocode for computing $v_{x}(\alpha)$ using a trained diffusion model. As in the case of $v_{\mathrm{FID}}(\alpha)$, this involves simulating the forward process on the training data and measuring changes in the predicted data across adjacent timesteps. Hyperparameters are summarized in Table 9. We use the same configuration for all our experiments, regardless of the diffusion type or dataset.